Calculate the formula unit masses of ZnO, Na2O, K2Co3 given atomic masses of Zn=65 u, Na = 23 u, K = 39 u, C = 12u, and O = 16 u.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The formula unit mass of
(i) ZnO = 65 u + 16 u =.81 u
(ii) Na2O = (23 u × 2) + .16 u : = 46u+ 161.u.
= 62 u
(iii) K2CO3 = (39 u × 2) + 12 u + 16 u × 3
= 75 u + 12u = 48u = 138u

Thanks